OpenStudy (anonymous):

HELP WILL MEDAL 2n-7 ---- = 15 3 n=__ What is n? x - - 8 = 12 4 x=__ What is x?

OpenStudy (gorv):

multiply both side by 3

OpenStudy (gorv):

then add 7 on both side

OpenStudy (gorv):

and after that divide both side by 2

OpenStudy (anonymous):

n = 11

OpenStudy (anonymous):

ok, what about the other question? :) n thx

OpenStudy (gorv):

N=11 is the wrong ans @Algebruhh

OpenStudy (anonymous):

plug 11 in for n. 2*11-7=15 . Not Wrong

OpenStudy (anonymous):

^^^

OpenStudy (gorv):

in denominator 3 is there

OpenStudy (anonymous):

n=11

OpenStudy (gorv):

u have to divide by 3 also

OpenStudy (gorv):

\[\frac{ 2n-7 }{ 3 }=15 \]

OpenStudy (gorv):

it is the problem

OpenStudy (gorv):

now n=11??

OpenStudy (gorv):

multiply both side by 3

OpenStudy (anonymous):

where does it say / by 3 ?

OpenStudy (gorv):

2n-7=15*3 2n-7=45

OpenStudy (anonymous):

so divide 45/3?

OpenStudy (gorv):

u r very slow

OpenStudy (gorv):

first we multiplied by 3 we got 45

OpenStudy (gorv):

@Algebruhh

OpenStudy (gorv):

2n-7=45 add 7 on each side 2n=45+7 2n=52 divide both side by 2

OpenStudy (gorv):

n=52/2

OpenStudy (anonymous):

so now I have 2n-14=52 then I divide 2

OpenStudy (gorv):

why u r so slow lol

OpenStudy (gorv):

2n=52 divide both side by 2

OpenStudy (gorv):

check out each step above u will get the ans

OpenStudy (gorv):

n=52/2=26 n=26

OpenStudy (anonymous):

srry ppl r diffent. have different pases. the answer is 26 correct? plus my internet laggs

OpenStudy (anonymous):

I got a 60 on it. btw it was 11 :/